# Ultra-Fast Language Generation via Discrete Diffusion Divergence Instruct

### v2

@@ Abstract
Sample text from 162M GPT-2 (1024 NFEs). Perplexity=$37.50$. It was hard, you know? Working with kids in school, holding a Pride event, creating, to me, a really awesome community for everyone to play a part in.
Sample text from 169M base model (1024 NFEs). Perplexity=$38.53$. I thought that when I was in high school, it would be too hard to talk to people who were like my classmates, but now I feel like I can talk to them better.
Sample text from 169M DiDi-Instruct (16 NFEs). Perplexity=$30.99$. The same thing is happening with classmates. As I get older, I think the best part is when I see people my age; I guess I get bored with that.
Figure 1: Perplexity vs. NFEs [^1] Baselines for comparison include GPT-2 Small, masked diffusion language models (MDLM; Sahoo et al. (2024)), diffusion duality (DUO; Sahoo et al. (2025)), and self-distillation through time (SDTT; Deschenaux and Gulcehre (2025))..
@@ 1 Introduction
Fast language sequence generation has been a long-standing goal for large-scale AI systems. Being one of the most typical large language models (LLMs), the Auto-regressive (AR) models have achieved remarkable success across a wide spectrum of natural language tasks (Brown et al., 2020; Meta AI, 2023; Achiam et al., 2023; Team et al., 2024). By predicting the next token from left to right using causal attention of Transformers (Vaswani et al., 2017), such models can scale up to billions or even trillions of parameters and deliver state‑of‑the‑art performance (Touvron et al., 2023). However, the interior mechanism that underpins their success imposes an inherent bottleneck: tokens are generated sequentially one at a time, limiting parallelism and throughput at scales (Leviathan et al., 2023; Kim et al., 2023; Ning et al., 2024). Even with advanced computational techniques, such as KV caches, AR models still have a significant throughput ceiling.
@@
Emerging as a competitive language generation paradigm, discrete diffusion large language models (dLLMs) (Li et al., 2022a) reinterpret text generation as an iterative text tokens denoising process inspired by continuous diffusion models for images (Song et al., 2021b; Ho et al., 2020). Starting from a fully corrupted token sequence, dLLMs first train and then apply a learned sequence denoiser several times to recover the clean text token sequences. As a result, this inference paradigm enables dLLMs to use the bidirectional attention inherited from the transformer network, and therefore, can generate sequences with fewer numbers of function evaluations (NFEs) than AR models.
@@
Despite their good performances and improved efficiency, dLLMs still face the inference efficiency bottleneck in generation steps. For instance, on the OpenWebText benchmark, dLLMs still need up to 256 NFEs to match the performance of a GPT2 model when generating a sequence of a length of 1024. To mitigate these issues, several distillation methods for dLLMs have been proposed to facilitate few-step generators to produce high-quality samples without performance degradation. For example, Deschenaux and Gulcehre (2025) proposed SDTT that trains a student to match a teacher’s predictions across time-steps, allowing the student to generate at least 32 tokens at once and achieving speedups up to eight times over AR models. Sahoo et al. (2025) introduced DUO that connects dLLMs to Gaussian diffusion and introduces curriculum learning and discrete consistency distillation to enable few‑step sampling. Zhu et al. (2025b) proposed the DiMO that matches the token‑level distribution of dLLMs and uses a token initialization strategy to produce a one‑step generator. We defer a detailed, side-by-side introduction and discussions of their distillation mechanisms in Appendix B.3. Though these early studies have pioneered the frontier of fast language generation, their generation throughput still has much room to improve: on OpenWebText, even the most recent SDTT can not match the performance of GPT2 baseline within 32 NFEs. Moreover, existing distillation methods are more straightforward in principle rather than well-established theoretical frameworks.
@@
In this work, we introduce Discrete Diffusion Divergence Instruct (DiDi-Instruct), a novel distillation framework for fast language generation. The principle of DiDi-Instruct is to minimize an Integral Kullback–Leibler (IKL) divergence introduced by the seminal work of Luo et al. (2023b) in the continuous diffusion model. By minimizing IKL between the distributions of a few-step student generation model and a pre-trained dLLM, the student can match the ability of dLLM in a distribution-matching manner with significantly improved efficiency. However, directly adapting IKL divergence to MDMs presents unique challenges due to discrete, non-differentiable operations that build the foundation of language sequence generation. To overcome these challenges, our key contributions are a comprehensive and effective solution through innovations in objective design, training stability, and inference efficiency. Specifically, we make the following contributions:
@@
  • Principled Training Method for Fast (Language) Sequence Generation: We reformulate the distillation objective from a general policy gradient perspective, deriving a simple yet tractable update rule for the few-step student according to some reward function. With an adversarial language discriminator to estimate the log-density ratio (reward) between the teacher dLLM and the student, we introduce a practical DiDi-Instruct algorithm that trains the few-step student, accompanied by an assistant discriminator.
@@
  • Simple yet Effective Techniques in Training and Inference: We introduce grouped reward normalization, intermediate-state matching, and the reward-guided ancestral sampler (RGAS) that significantly improve the training stability, the model coverage, and the inference performances, reducing the generation perplexity by 30%.
@@
  • State-of-the-art Fast Sequence Generation: DiDi-Instruct achieves new state-of-the-art performance on the OpenWebText benchmark: consistently lower PPL across 8 to 128 NFEs (Figure 1), negligible entropy loss, and over $20\times$ faster distillation; detailed ablations, model scaling, and protein sequence generation further confirm its robustness.
@@
To the best of our knowledge, DiDi-Instruct is the first framework to successfully apply distribution-matching distillation to MDM-based text generation, along with record-breaking performances in few-step language sequence generation.
2 Related Work
Continuous-space diffusion distillation. Diffusion models (Song et al., 2021b; Ho et al., 2020; Sohl-Dickstein et al., 2015) added Gaussian noises to data with a forward diffusion process and then learn the backward stochastic differential equations (SDEs) with so-called score neural networks. Early acceleration reduced the steps of iterative samplers using training-free high-order ODE solvers (Song et al., 2021a; Lu et al., 2022; Karras et al., 2022; Xue et al., 2023). However, solver-based acceleration suffered from discretization error in less than 10 generation steps. Luhman and Luhman (2021) and Salimans and Ho (2022) first introduced training the few-step student models by learning the consecutive trajectory mapping of diffusion solvers. The seminal work of Song et al. (2023) and Luo et al. (2023b) opens the one-step generation of diffusion models through different distillation principles. Specifically, consistency models (Song et al., 2023) distill the few-step models with the trajectory consistency principle, resulting in strong performances (Song and Dhariwal, 2024; Lu and Song, 2025; Geng et al., 2025b; Luo et al., 2023a; Kim et al., 2024; Geng et al., 2025a). Diff-Instruct (Luo et al., 2023b) introduces the distribution matching principle that distills one-step generative models, resulting in leading efficient image generative models (Wang et al., 2025; Luo et al., 2025a, 2024; Luo, 2025; Xu et al., 2025; Yin et al., 2024b, a; Xie et al., 2024; Fan et al., 2023; Zhou et al., 2024; Huang et al., 2024; Luo et al., 2025b, c). It is worth noting that, in this paper, we are inspired by the distribution-matching principle introduced by Diff-Instruct, but in the context of challenging discrete language sequence generation.
Discrete diffusion for language. Discrete modeling began with categorical transitions via argmax flows (Hoogeboom et al., 2021). Later, Austin et al. (2021) introduced structured transition matrices and auxiliary cross-entropy training. For language modeling, two principal families of dLLMs have emerged. Masked diffusion models (MDMs) (Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2025) treat the forward process as progressive masking of tokens. At the highest noise level, every token is replaced by a special [MASK] symbol; the reverse process gradually unmask tokens. Once a token is unmasked, it remains fixed, but many tokens can be denoised simultaneously. Intuitively, MDMs essentially train BERT (Devlin et al., 2019) under a hierarchy of noise levels, motivated by a scaling-based rationale for generation (He et al., 2023; Sahoo et al., 2024). Uniform‑state diffusion models (USDMs) (Austin et al., 2021; Gulrajani and Hashimoto, 2023; Schiff et al., 2025; Sahoo et al., 2025) instead corrupt tokens by sampling from a uniform distribution over the vocabulary. Consequently, each token can change multiple times during sampling, enabling self‑correction and strong theoretical connections to continuous Gaussian diffusion. Subsequent dLLM work both elucidates training and decoding challenges (Zheng et al., 2025; Kim et al., 2025) and scales models to the billion-parameter range (Nie et al., 2025; Zhu et al., 2025a; Song et al., 2025). Our distillation framework, which adopts ideas from continuous-time IKL, is specifically designed for dLLMs to ensure tractable training and inference.
@@ 3 Preliminary
We first introduce the notations of MDMs (Section 3.1) and the integral KL divergence (Section 3.2), for the preparation of the DiDi-Instruct algorithm in Section 4. We denote a vocabulary of size $K$. A one-hot vector represents each element in this vocabulary. The set of all such vectors is defined as $\mathcal{V}=\big\{\mathbf{x}\in\{0,1\}^{K}\mid\sum_{j=1}^{K}x_{j}=1\big\}.$ Within this vocabulary, the $K$-th token is reserved for the special [MASK] token. Its one-hot encoding is the standard basis vector $\mathbf{m}$, where the $K$-th element is $1$ and all others are 0. To generate a sequence $\mathbf{x}^{1:L}=(\mathbf{x}^{1},\dots,\mathbf{x}^{L})\in\mathcal{V}^{L}$ of length $L$, we assume that the process factorizes across positions, i.e., each index $\ell\in\{1,\dots,L\}$ evolves independently. We adopt a continuous time index $t\in[0,1]$ with a monotonically decreasing noise schedule $\alpha_{t}\in[0,1]$ satisfying $\alpha_{0}\approx 1$ and $\alpha_{1}\approx 0$. Moreover, we denote $K$-simplex as $\Delta^{K}$.
@@ 3.1 Masked Diffusion Models
Forward process. In the forward process, once a token transitions to the masked state $\mathbf{m}$, it never reverts. With independent per-token transitions, $\mathbf{x}^{1:L}$ converges to fully masked sequence $(\mathbf{m},\dots,\mathbf{m})$ with probability $1$ as $t\to 1$ (Sahoo et al., 2024). In the forward corruption process, the clean token $\mathbf{x}\in\mathcal{V}$ transitions to a latent state $\mathbf{z}_{t}\in\mathcal{V}$ at time $t\in(0,1]$. This absorbing state process ensures that $\mathbf{z}_{t}$ either remains the original token $\mathbf{x}$ with probability $\alpha_{t}$ or transitions into $\mathbf{m}$ with probability $1-\alpha_{t}$. The forward corruption kernel is formalized as:
@@
where $\mathrm{Cat}\big(\mathbf{z};\bm{\pi})$ denotes a Categorical distribution over the one-hot vector $\mathbf{z}\in\mathcal{V}$ with $\bm{\pi}\in\Delta^{K}$. For $s<t$, the exact posterior takes two cases:
@@
Intuitively, an unmasked token implies a deterministic past, whereas a masked token implies a past probabilistic mixture governed by the noise schedule.
Backward process. The reverse process of MDMs iteratively reconstructs the original uncorrupted sequence from the noisy input by replacing masked tokens. This is achieved through a learned neural network $\mathbf{p}_{\theta}:\mathcal{V}\times[0,1]\rightarrow\Delta^{K}$ that predicts the clean token $\mathbf{x}$ from its corrupted version $\mathbf{z}_{t}$ at time $t$. The output of $\mathbf{p}_{\theta}(\mathbf{z}_{t},t)$ satisfies that (i) the predicted distribution forms valid class probabilities that sum to $1$, (ii) the prediction $\mathbf{p}_{\theta}(\mathbf{z}_{t},t)$ assigns zero probability mass to the [MASK] token, and (iii) once a token is unmasked, its state remains unchanged in subsequent steps. When applied to sequences $\mathbf{x}^{1:L}$, the denoising process is assumed to be token-wise independent. A network $\mathbf{p}_{\theta}^{1:L}(\mathbf{z}_{t}^{1:L},t)$ predicts the distribution for the entire sequence, where $\mathbf{p}_{\theta}^{\ell}(\mathbf{z}_{t}^{1:L},t)$ specifies the distribution for the $\ell$-th token. The reverse transition from $\mathbf{z}_{t}$ to $\mathbf{z}_{s}$ is defined by $\mathcal{P}_{\theta}(\mathbf{z}_{s}\mid\mathbf{z}_{t})=\mathcal{Q}(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}=\mathbf{p}_{\theta}(\mathbf{z}_{t},t))$. We extend this to sequences by factorizing the process over the sequence length $L$: $\mathcal{P}_{\theta}(\mathbf{z}_{s}^{1{:}L}\mid\mathbf{z}_{t}^{1{:}L})=\prod_{\ell=1}^{L}\mathcal{Q}(\mathbf{z}^{\ell}_{s}\mid\mathbf{z}^{\ell}_{t},\mathbf{p}_{\theta}^{\ell}(\mathbf{z}_{t}^{1{:}L},t))$ In continuous time, minimizing the negative evidence lower bound (NELBO) provides a tractable training objective (Shi et al., 2024):
$\mathcal{L}^{\infty^{L}}_{\mathrm{NELBO}}(\theta)=\mathbb{E}_{q}\left[\int_{0}^{1}\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}\frac{1}{1-\alpha_{t}}\Big[\sum_{\ell:\mathbf{z}_{t}^{\ell}=\mathbf{m}}\mathbf{x}^{\ell^{\top}}\log\mathbf{p}_{\theta}^{\ell}(\mathbf{z}_{t}^{1{:}L},t)\Big]\mathrm{d}t\right].$ (3)
@@ 3.2 Integral KL divergence
The iterative reverse process in MDMs necessitates a large NFEs, resulting in high inference latency. To mitigate this, we frame accelerated generation as a distillation task, where a multi-step teacher guides a student model to produce high-quality samples in significantly fewer steps. Since the student must learn the teacher’s entire generative trajectory, we consider the following distillation framework that aligns their marginal distributions across the full time continuum.
Let $\mathbf{q}_{\theta}(\mathbf{z_{t}},t)$ denote the forward marginal and $\mathbf{q}_{\nu}(\mathbf{z_{t}},t)$ the student marginal parameterized by $\nu$ [^1] Unless otherwise specified, $\mathbf{q}_{\theta}$ and $\mathbf{q}_{\nu}$ denote the full collections $\mathbf{q}_{\theta}^{1:L}(\mathbf{z_{t}},t)$ and $\mathbf{q}_{\nu}^{1:L}(\mathbf{z_{t}},t)$.. To distill knowledge from $\mathbf{q}_{\theta}$ to $\mathbf{q}_{\nu}$ across different noise levels $t\in[0,1]$, we minimize the IKL divergence (introduced by Luo et al. (2023b)), which aggregates the discrepancy between $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$:

### v3

@@ Abstract
Sampled text from 162M GPT-2 (1024 NFEs). Generative Perplexity=$37.50$. It was hard, you know? Working with kids in school, holding a Pride event, creating, to me, a really awesome community for everyone to play a part in.
Sampled text from 169M base model (1024 NFEs). Generative Perplexity=$38.53$. I thought that when I was in high school, it would be too hard to talk to people who were like my classmates, but now I feel like I can talk to them better.
Sampled text from 169M DiDi-Instruct (16 NFEs). Generative Perplexity=$30.99$. The same thing is happening with classmates. As I get older, I think the best part is when I see people my age; I guess I get bored with that.
Figure 1: Perplexity vs. NFEs [^1] Baselines include GPT-2, masked diffusion language models (MDLM; Sahoo et al. (2024)), diffusion duality (DUO; Sahoo et al. (2025)), and self-distillation through time (SDTT; Deschenaux and Gulcehre (2025))..
@@ 1 Introduction
Fast language sequence generation has been a long-standing goal for large-scale AI systems. Auto-regressive (AR) large language models (LLMs) have achieved remarkable success across a wide range of natural language tasks (Brown et al., 2020; Meta AI, 2023; Achiam et al., 2023; Team et al., 2024). By predicting the next token from left to right using causal attention of Transformers (Vaswani et al., 2017), AR models can scale to billions or even trillions of parameters while delivering state‑of‑the‑art performance (Touvron et al., 2023). However, the interior mechanism that enables this success also imposes an inherent bottleneck: tokens must be generated sequentially, which limits parallelism and constrains throughput at scales (Leviathan et al., 2023; Kim et al., 2023; Ning et al., 2024). Even with recent computation optimization, such as KV caches, AR models still face a significant throughput ceiling.
@@
Inspired by continuous diffusion models for images (Song et al., 2021b; Ho et al., 2020), diffusion large language models (dLLMs) (Li et al., 2022) reinterpret text generation as an iterative denoising process over token sequences, offering a competitive alternative to the AR model. Concretely, a dLLM starts from a fully corrupted sequence and learns to progressively denoise it into a clean text sequence. This inference paradigm leverages bidirectional attention in the Transformer, allowing dLLMs generate sequences with fewer numbers of function evaluations (NFEs) than AR models.
@@
Despite their strong performance and improved efficiency, dLLMs still require many denoising steps during inference, creating an inference time bottleneck. For example, on the OpenWebText benchmark, dLLMs still need up to 256 NFEs to match the performance of a GPT2 model when generating sequences of length 1024. To mitigate this limitation, several distillation methods for dLLMs have been proposed to train few-step generators to produce high-quality samples without performance degradation. Notably, Deschenaux and Gulcehre (2025) trains a student to match a teacher’s predictions across time-steps, enabling the student to generate at least 32 tokens per step and achieving up to $8\times$ speedups over AR models. Later, Sahoo et al. (2025) relates dLLMs to Gaussian diffusion and introduces curriculum learning and discrete consistency distillation to support few‑step sampling. Zhu et al. (2025b) matches the token‑level distribution of dLLMs and uses a token initialization strategy to obtain a one‑step generator. We defer a detailed, side-by-side introduction and discussions of their distillation mechanisms to Appendix A.3. While these studies advance the frontier of fast language generation, their generation throughput still leaves substantial room for improvement: on OpenWebText, even the most recent SDTT can not match the performance of GPT2 baseline within 32 NFEs. Moreover, existing distillation methods are motivated by heuristic design choices, lacking well-established theoretical justification.
@@
In this work, we introduce Discrete Diffusion Divergence Instruct (DiDi-Instruct), a novel distillation framework for fast language generation. The learning objective of DiDi-Instruct is to minimize the Integral Kullback–Leibler (IKL) divergence, introduced by the seminal work of Luo et al. (2023b) for the continuous diffusion model. By minimizing IKL between the distributions of a few-step student generator and a pre-trained teacher dLLM, the student is trained to match the teacher’s generation distribution, while achieving substantially improved inference efficiency. Directly adapting IKL to masked diffusion models (MDMs), however, is nontrivial due to the discrete and non-differentiable operations inherent in dLLM generation. To overcome these challenges, we develop a comprehensive solution spanning objective design, training stability, and inference efficiency. Specifically, we make the following contributions:
@@
  • Principled Training for Fast Sequence Generation: We reformulate the distillation objective from a general policy gradient perspective, deriving a simple yet tractable update rule for the few-step student according to some reward function. Using an adversarial language discriminator to estimate the log-density ratio (reward) between the teacher dLLM and the student, we introduce a practical DiDi-Instruct algorithm that trains the few-step student, paired with an assistant discriminator.
@@
  • Simple yet Effective Techniques in Training and Inference: We introduce grouped reward normalization, intermediate-state matching, and the reward-guided ancestral sampler (RGAS) that substantially improve the training stability, the model coverage, and the inference performances, reducing the perplexity of generated sequences by 30%.
@@
  • State-of-the-Art Fast Sequence Generation: DiDi-Instruct achieves new state-of-the-art performance on the OpenWebText benchmark: consistently lower PPL across 8 to 128 NFEs (Figure 1), negligible entropy loss, and over $20\times$ faster distillation; detailed ablations, model scaling, downstream tasks, and protein sequence generation further confirm its robustness.
@@
To the best of our knowledge, DiDi-Instruct is the first framework to successfully apply distribution-matching distillation to MDM-based text generation, along with record-breaking performances in few-step language sequence generation. A discussion of related work is deferred to Appendix A.
@@ 2 Preliminary
This section introduces the notations for MDMs (Sec. 2.1) and the integral KL divergence (Sec. 2.2), for preparation of the DiDi-Instruct algorithm in Section 3. We consider a vocabulary of size $K$, where each token in the vocabulary is represented by a one-hot vector. The set of all such vectors is $\mathcal{V}=\{\mathbf{x}\in\{0,1\}^{K}|\sum_{j=1}^{K}x_{j}=1\}$. The $K$-th token is reserved for the special [MASK] token, whose one-hot encoding is noted by vector $\mathbf{m}$ with the $K$-th entry being $1$ and all others being 0. To generate a length-$L$ sequence $\mathbf{x}^{1:L}=(\mathbf{x}^{1},\dots,\mathbf{x}^{L})\in\mathcal{V}^{L}$, we assume the process factorizes across positions, i.e., each index $\ell\in\{1,\dots,L\}$ evolves independently. We adopt a continuous time index $t\in[0,1]$ with a monotonically decreasing noise schedule $\alpha_{t}\in[0,1]$ satisfying $\alpha_{0}\approx 1$ and $\alpha_{1}\approx 0$. Finally, we denote $K$-simplex as $\Delta^{K}$.
@@ 2.1 Masked Diffusion Models
Forward process. In the forward corruption process, once a token transitions to the masked state $\mathbf{m}$, it never reverts. Under independent per-token transitions, $\mathbf{x}^{1:L}$ converges to the fully masked sequence $(\mathbf{m},\dots,\mathbf{m})$ with probability $1$ as $t\to 1$ (Sahoo et al., 2024). Concretely, a clean token $\mathbf{x}\in\mathcal{V}$ transitions to a latent state $\mathbf{z}_{t}\in\mathcal{V}$ at time $t\in(0,1]$. This absorbing-state process implies that $\mathbf{z}_{t}$ either stays as the original token $\mathbf{x}$ with probability $\alpha_{t}$ or transitions to $\mathbf{m}$ with probability $1-\alpha_{t}$. The forward corruption kernel is:
@@
where $\mathrm{Cat}\big(\mathbf{z};\bm{\pi})$ denotes a Categorical distribution over the one-hot vector $\mathbf{z}\in\mathcal{V}$ with probability vector $\bm{\pi}\in\Delta^{K}$. For $s<t$, the exact posterior has two cases:
@@
The intuition of Equation (2) is: at time $t$, an unmasked token implies a deterministic pash, whereas a masked token considers a probabilistic mixture governed by the noise schedule.
Backward process. The reverse process of MDMs iteratively reconstructs the original uncorrupted sequence from the noised input by replacing masked tokens. This is achieved through a learned neural network $\mathbf{p}_{\theta}:\mathcal{V}\times[0,1]\rightarrow\Delta^{K}$ that predicts the clean token $\mathbf{x}$ from its corrupted version $\mathbf{z}_{t}$ at time $t$. The output of $\mathbf{p}_{\theta}(\mathbf{z}_{t},t)$ satisfies that (i) the predicted distribution forms valid categorical distribution that sum to $1$, (ii) the prediction $\mathbf{p}_{\theta}(\mathbf{z}_{t},t)$ assigns zero probability mass to the [MASK] token, and (iii) once a token is unmasked, its state remains fixed in all subsequent steps.
When applied to sequences $\mathbf{x}^{1:L}$, the denoising process is assumed to be token-wise independent. A network $\mathbf{p}_{\theta}^{1:L}(\mathbf{z}_{t}^{1:L},t)$ predicts the distribution over the entire sequence, where $\mathbf{p}_{\theta}^{\ell}(\mathbf{z}_{t}^{1:L},t)$ specifies the distribution for the $\ell$-th token. The reverse transition from $\mathbf{z}_{t}$ to $\mathbf{z}_{s}$ is defined as $\mathcal{P}_{\theta}(\mathbf{z}_{s}\mid\mathbf{z}_{t})=\mathcal{Q}(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}=\mathbf{p}_{\theta}(\mathbf{z}_{t},t))$. We extend this to sequences by factorizing the process over the sequence of length $L$: $\mathcal{P}_{\theta}(\mathbf{z}_{s}^{1{:}L}\mid\mathbf{z}_{t}^{1{:}L})=\prod_{\ell=1}^{L}\mathcal{Q}(\mathbf{z}^{\ell}_{s}\mid\mathbf{z}^{\ell}_{t},\mathbf{p}_{\theta}^{\ell}(\mathbf{z}_{t}^{1{:}L},t))$ In continuous time, minimizing the negative evidence lower bound (NELBO) provides a tractable training objective (Shi et al., 2024):
$\mathcal{L}^{\infty^{L}}_{\mathrm{NELBO}}(\theta)=\mathbb{E}_{q}\left[\int_{0}^{1}\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}\frac{1}{1-\alpha_{t}}\left[\sum_{\ell:\mathbf{z}_{t}^{\ell}=\mathbf{m}}\mathbf{x}^{\ell^{\top}}\log\mathbf{p}_{\theta}^{\ell}(\mathbf{z}_{t}^{1{:}L},t)\right]\mathrm{d}t\right].$ (3)
@@ 2.2 Integral KL divergence
Let $\mathbf{q}_{\theta}(\mathbf{z_{t}},t)$ denote the forward teacher marginal and $\mathbf{q}_{\nu}(\mathbf{z_{t}},t)$ the student marginal parameterized by $\nu$ [^1] Unless otherwise specified, $\mathbf{q}_{\theta}$ and $\mathbf{q}_{\nu}$ denote the full collections $\mathbf{q}_{\theta}^{1:L}(\mathbf{z_{t}},t)$ and $\mathbf{q}_{\nu}^{1:L}(\mathbf{z_{t}},t)$.. To distill knowledge from $\mathbf{q}_{\theta}$ to $\mathbf{q}_{\nu}$ across different noise levels $t\in[0,1]$, we minimize the integral Kullback-Leibler (IKL) divergence (introduced in Luo et al. (2023b)), which aggregates the discrepancy between $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$ over the time domain: